OpenStudy (anonymous):

Y=2x + radical 9 - 35x find the x and y intercepts

OpenStudy (texaschic101):

y = 2x + sq rt 9 - 35x y = 2x + 3 - 35x y = -33x + 3 y intercept is 3 ===> y intercept (0,3) x intercept can be found by subbing in 0 for y : 0 = - 033x + 3 33x = 3 x = 11 ===> x intercept (11,0)

OpenStudy (anonymous):

Wouldn't x be 1/3,0